find an equation of the line that has a y-intercept of 2 that is parallel to the graph of the line 4x+2y=8
Answer by Edwin McCravy(20086)

find an equation of the line that has a y-intercept of 2 that is parallel to the graph of the line 4x+2y=8
4x+2y=8

can be divided through by 2

2x+y=4

Parallel lines only differ by the constant terms. So we drop off the 4

2x+y=?

Substitute in the y-intercept which is the point (0,2)

2(0)+2=?

   0+2=2

So the answer is

2x+y=2

Edwin